OpenStudy (anonymous):

am i right? What value of k makes the statement true? –(k + 12) = 6 + (–12) A. –18 B. –6 C. 6 D. 18 my answer is:b

OpenStudy (igreen):

\(–(k + 12) = 6 + (–12)\) Distribute -1 into the parenthesis: \(-k - 12 = 6 - 12\) Subtract: \(-k - 12 = -6\) Add 12 to both sides: \(-k = 6\) Multiply -1 to both sides: \(k = -6\)

OpenStudy (igreen):

You're correct!

OpenStudy (anonymous):

kk